A human red blood cell has on average a diameter of 8.0E+0 microns (1 micron = 1ųm). What is this in nanometers ?
1
Expert's answer
2020-09-23T06:16:48-0400

A human red blood cell has on average a diameter of 8.0E+0 microns (1 micron = 1ųm)


( 8.0E + 0 ) * 10 -6 m = ( 8000 .0 E + 0 ) * 10 -9 m = ( 8000.0 E + 0 ) nm .